OpenStudy (anonymous):

What is the simplified form of each expression? 7^4 A. 343 B. 16,384 C. 16,807 D. 2,401

OpenStudy (anonymous):

D. 2,401

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 7^4 \Longrightarrow 7 * 7 * 7 * 7 = ? }\)

OpenStudy (anonymous):

7 * 7 * 7 * 7 = is what

OpenStudy (apoorvk):

\[a^n = a*a*a*....*a ----> n times\] so you gave to multiply 7 by itself 4 times, to find out 7^4

Parth (parthkohli):

@swiftskier96 I respect you spending time on the question and thinking, but direct answering is against the Code Of Conduct on OpenStudy.

OpenStudy (anonymous):

7*7=49 49*7= ...

OpenStudy (anonymous):

7 * 7 = 49 49 * 7 = 343 343 * 7 = 2, 401 is the answer